# MIM-Refiner: A Contrastive Learning Boost from Intermediate Pre-Trained Representations

### v2

@@ Abstract
We introduce MIM (Masked Image Modeling)-Refiner, a contrastive learning boost for pre-trained MIM models. MIM-Refiner is motivated by the insight that strong representations within MIM models generally reside in intermediate layers. Accordingly, MIM-Refiner leverages multiple contrastive heads that are connected to different intermediate layers. In each head, a modified nearest neighbor objective constructs semantic clusters that capture semantic information which improves performance on downstream tasks, including off-the-shelf and fine-tuning settings.
The refinement process is short and simple – yet highly effective. Within a few epochs, we refine the features of MIM models from subpar to state-of-the-art, off-the-shelf features. Refining a ViT-H, pre-trained with data2vec 2.0 on ImageNet-1K, sets a new state-of-the-art in linear probing (84.7%) and low-shot classification among models that are pre-trained on ImageNet-1K. At ImageNet-1K 1-shot classification, MIM-Refiner advances the state-of-the-art to 64.2%, outperforming larger models that were trained on up to 2000 times more data such as DINOv2-g, OpenCLIP-G and MAWS-6.5B. https://ml-jku.github.io/MIM-Refiner
@@ 1 Introduction
Self-supervised learning (SSL) has attracted considerable attention, owing to its data efficiency and generalization ability [49]. SSL leverages pre-taining tasks and creates intricate input representations without the need for explicit supervision via expensive annotations. In computer vision, Masked Image Modeling (MIM) [12, 69, 58] has emerged as one of the prevalent SSL pre-training paradigms, enabling an efficient pretraining of large models on unlabeled data by reconstructing masked parts of the input images.
@@
In this work, we first analyze pre-trained MIM models, where we find that MIM models have different types of blocks: those that mainly improve the pre-training objective and others that are responsible for abstraction within the MIM encoder. The origin of this behavior can be traced back to the fact that MIM architectures usually comprise a large ViT encoder together with a very light-weight decoder. For larger models, the light-weight decoder reaches a point, where it cannot further improve the pre-training objective on its own and passes part of the reconstruction task back to the last encoder blocks. Consequently, the feature quality for downstream tasks of the later blocks degrades, and, somewhat unusual, the representation quality peaks in the middle blocks of the encoder.
@@
We introduce Nearest Neighbor Alignment (NNA) as a new ID variant. Given a batch of samples, NNA aligns each sample with its NN and repels each sample from all other samples in the batch.
Experimentally, we show that within few epochs, MIM-Refiner refines the features of MIM models from subpar to state-of-the-art, off-the-shelf features, see Figure 2. For example, a refined ViT-H, pre-trained with data2vec 2.0 [6] on ImageNet-1K [20], achieves new state-of-the-art results in linear probing (84.7%, Figure 1), $k$-NN accuracy (82.3%), $k$-means cluster accuracy (67.3%) and low-shot classification among models pre-trained on ImageNet-1K. In ImageNet-1K 1-shot classification it sets a new state-of-the-art of 64.2%, outperforming larger models that were trained on up to 2000x more data such as DINOv2-g, OpenCLIP-G and MAWS-6.5B.
@@
    We analyze MIM models, and find that MIM models have different types of blocks: those that mainly improve the pre-training objective and others that are responsible for abstraction within the MIM encoder.
@@
    We introduce Nearest Neighbor Alignment (NNA), an ID objective that aligns each sample with its nearest neighbor while repelling other samples in a batch.
2 Challenges of Masked Image Modeling
@@
Figure 3: (a) MIM models are asymmetrically designed where the encoder has most of the parameters, as indicated by the percentages in the bars. (b) The representation quality of the encoder (measured by $k$-NN accuracy) peaks in the middle blocks before degrading when the later blocks take over parts of the decoder’s task. Various MIM models show such a pattern (see Appendix C.1) (c) The decline in representation quality in later blocks primarily contributes to the degradation in downstream performance. ID methods (represented by iBOT [81]) and our MAE-Refined do not suffer from this issue. (d) Correlation of the relative improvement of the reconstruction loss and the relative improvement in the $k$-NN accuracy per block. The relative $k$-NN accuracy improvement is obtained from the $k$-NN accuracy shown in (b). Similarly, the relative improvement of the reconstruction loss is obtained from Figure 8 in the appendix. Middle blocks form abstract representations (large improvements in the $k$-NN accuracy, almost no improvement in reconstruction loss), later blocks take over parts of the reconstruction task (decrease in the $k$-NN accuracy, large improvement in the reconstruction loss). Additional details can be found in Appendix E.2.
MIM models, such as MAE [29] and data2vec 2.0 [6], enable an efficient pre-training of large models. In terms of architecture, the encoder and decoder are intentionally designed asymmetrically. The encoder, on the one hand, is a large ViT, where discarding 75% of the input patches drastically reduces the cost of a forward pass. The decoder, on the other hand, operates on the full sequence length – by concatenating mask tokens to the encoded visible patches – and, thus, is typically very lightweight to compensate for the increased number of tokens (Figure 3(a)). As models increase in size, the decoder eventually reaches a point where it cannot further improve the pre-training objective on its own. Consequently, it begins to delegate a portion of the reconstruction task back to the last encoder blocks. This transfer adversely affects the feature quality for downstream tasks associated with those blocks (Figure 3(b)), especially when only few labels are available (Figure 3(c)). We observe this phenomenon by correlating the relative improvement in the reconstruction loss vs the $k$-NN accuracy (Figure 3(d)). Roughly speaking, the blocks of the encoder operate in three different regimes:
@@
When naïvely using the features of the last encoder block, those features are suited for reconstruction, but not particularly for downstream tasks. If lots of labels are available, this can be compensated by fine-tuning the last encoder blocks on the downstream task. However, if not enough labels are available, or the last encoder blocks are not fine-tuned, downstream performance suffers.
One would think that simply using a larger decoder solves these problems. However, there are multiple problems with this solution. (i) The decoder commonly operates on the full sequence length, making it costly to increase its size. (ii) Scaling the decoder can decrease performance as shown, for example, in MAE [29] (Table 1). (iii) Models that can use a deeper decoder (such as CrossMAE [27]) also show degrading representation quality in later blocks, as shown in Appendix Figure 6(c).
@@ 3 Method
Inspired by NN based contrastive learning [24, 4], we propose Nearest Neighbor Alignment (NNA). NN contrastive objectives introduce an inter-sample correlation by retrieving NNs of samples in a batch and subsequently applying an objective between the samples and their NNs. In practice, the NN retrieval is typically done by tracking samples from previous iterations in a first-in-first-out queue [24, 30]. Therefore, the samples in the queue are from a previous model state which creates a tradeoff between the benefit of the NN-swap and the worse signal from the out-of-sync samples. We argue that the NN-swap does not offer a benefit for negative samples, since there are already different images in every batch, and instead, degrades the signal from the contrastive objective. We therefore propose to use the NN only for the alignment of positive samples, as visualized on the right side of Figure 4. Omitting the NN-swap for the negatives gives a small but consistent improvement as ablated in Table 4(e). NNA is a variant of NNCLR and we visualize their difference in Figure 10.
Formally, given a batch of features $\mathcal{Z}=\{z_{1},\dots,z_{N}\}$ and a queue $\mathcal{Q}$ the NNA objective is:
@@
4 Related Work
4.1 Pre-training in Computer Vision
4.2 Combining MIM and ID
@@ 5 Experiments
We refine a series of MIM models, namely MAE [29, 64], data2vec 2.0 [6] (abbreviated as D2V2), dBOT [50] and CrossMAE [27]. These models were pre-trained on ImageNet-1K, which we also use for the refinement. For visual clarity, we compare our best model against previous state-of-the-art models and show the full result tables in the Appendix. We evaluate our models in classification (low- and many-shot), feature evaluation, clustering, transfer learning and semantic segmentation.
@@ 5.1 Ablation Study
We first investigate various design choices of our method in Table 4(f).
(a) Adding multiple heads at intermediate features improves $k$-NN accuracy by 0.8%. The best settings include the blocks with the best representation quality (see Figure 3(b)). Adding additional heads before the best blocks does not improve performance while increasing training costs. (d) The benefit doubles to 1.6% for a ViT-H since deeper models degrade more (see Figure 3(b)). Additionally, we investigate where to attach the ID heads in Appendix C.2 where we find that simply attaching ID head to the last 8 (or 12 for ViT-H) blocks to perform best across model scales.
(b) Freezing early blocks as a form of regularization to preserve MIM features (similar to related works [47, 40]) is not necessary. Note that we still use a layer-wise learning rate decay [17].
(c) Scheduling the loss weight of each intermediate head is not necessary, i.e.  simply summing all losses is sufficient to achieve good performances. “Uniform Decay” decays the loss weight of all intermediate heads during training. “Staggered Decay” starts by decaying the first head and gradually starts to decay more and more heads as training progresses. “Staggered Step” disables the first head after some time followed by gradually disabling more and more heads. “One Hot” trains the encoder with only one intermediate head at a time where training starts with the earliest intermediate head and gradually iterates through the heads. Details to the loss schedules are in Appendix E.13.
(e) Using the NN swap only for aligning the positive with the anchor is slightly better than also swapping the negatives with their NN.
(f) Relying only on the data-driven augmentation of the NN-swap [24] by omitting color/blur augmentations, is beneficial for certain downstream tasks such as extreme low-shot classification [47]. We show more results without color/blur augmentations in Appendix C.3.
Table 1: Ablation study by refining data2vec 2.0 ViT-L/16 unless noted otherwise. Default settings
@@ 5.2 Low-shot and Feature Evaluations
We evaluate the ability of our models to perform low-shot classification in Table 2. Additionally, linear probing and $k$-NN accuracy are reported which are computed from the features of the frozen encoder. These metrics are typically correlated to low-shot performance as it indicates that the representation is already linear separable which eases drawing decision boundaries given only few labels. For linear probing, we use the protocol of DINOv2 [56] which includes using features of the last four blocks in its hyperparameter grid. Therefore, linear probing evaluates the feature representation at the end of the encoder, while $k$-NN accuracy evaluates only the features of the last block.
MIM-Refiner drastically improves upon MIM models and other SSL models. In the 1-shot settings D2V2-Refined-H sets a new state-of-the-art of 64.2%, outperforming the 63.6% of MAWS-6.5B [64] which is pre-trained on a private dataset with approximately 2000 times the size of ImageNet-1K.
@@ 5.3 Cluster Evaluations
Table 3 shows that MIM-Refiner greatly improves various clustering metrics. The reached ACC of 67.4% outperforms the current state-of-the-art of 61.6% reached by TEMI MSN [1]. Figure 5 illustrate this drastic increase in cluster performance and class separation visually.
@@ 5.4 Transfer Learning Evaluations
Table 4: Transferring MIM-Refiner models to other datasets. “VTAB-6” reports the average accuracy over six datasets from the VTAB benchmark [78] and “VTAB-1K” is the average over all 19 datasets of the VTAB-1K benchmark. ADE20K reports the mean intersection over union (mIoU) of a semantic segmentation probe. MIM-Refiner learns general features that can easily be transferred to various datasets and tasks. Table 11 confirms this finding on additional models and Appendix C.7 shows individual VTAB performances for VTAB-6 probing and VTAB-1K fine-tuning.
@@ 5.5 Fine-tuning with Large Amounts of Labels
MIM models typically outperform ID methods when given enough labels to fine-tune the model. As our refinement process employs an ID objective, we investigate whether MIM-Refiner involuntarily degrades performance given an abundance of labels. To this end, we fine-tune MIM models and their refined version on ImageNet-1K [20], iNat18 [35] and ADE20K [80] using a linear classification/segmentation head. Table 5 shows a small but consistent improvent of MIM-Refiner models.
Table 5: Full fine-tuning using 100% of the labels. MIM-Refiner consistently improves performance slightly even though ID methods typically perform worse than MIM models on this benchmark. Appendix Table 14 confirms this pattern on additional MIM models. “IN-Rob.” shows the average performance of the trained IN-1K classifier on robustness datasets (individual results in Table 15).
@@
Model IN-1K IN-Rob. iNat18 ADE20K IN-1K IN-Rob. iNat18
MAE [29] 85.7 53.9 80.7 47.6 86.7 60.3 82.7
MAE-Ref. 85.6 54.4 80.9 48.3 86.9 60.9 83.3
dBOT [50] 85.8 55.3 81.9 51.0 87.1 63.7 84.1
dBOT-Ref. 85.9 55.3 82.1 51.1 87.1 64.0 84.5
Mugs [82] 85.2 46.4 76.9 48.5 - -

### v4

# MIM-Refiner: A Contrastive Learning Boost from Intermediate Pre-Trained Masked Image Modeling Representations

@@ Abstract
We introduce MIM (Masked Image Modeling)-Refiner, a contrastive learning boost for pre-trained MIM models. MIM-Refiner is motivated by the insight that strong representations within MIM models generally reside in intermediate layers. Accordingly, MIM-Refiner leverages multiple instance discrimination (ID) heads that are connected to different intermediate layers. In each head, a nearest neighbor ID objective constructs clusters that capture semantic information which improves performance on downstream tasks, including off-the-shelf and fine-tuning settings.
The refinement process is short and simple – yet highly effective. Within a few epochs, we refine the features of MIM models from subpar to state-of-the-art, off-the-shelf features. Refining a ViT-H, pre-trained with data2vec 2.0 on ImageNet-1K, sets a new state-of-the-art in linear probing (84.7%) and low-shot classification among models that are pre-trained on ImageNet-1K. MIM-Refiner efficiently combines the advantages of MIM and ID objectives, enabling scaling ID objectives to billion parameter models using relatively little compute. MIM-Refiner compares favorably against previous state-of-the-art SSL models on various benchmarks such as low-shot classification, long-tailed classification and semantic segmentation.
@@ 1 Introduction
Self-supervised learning (SSL) has attracted considerable attention, owing to its data efficiency and generalization ability [54]. SSL leverages pre-training tasks and creates intricate input representations without the need for explicit supervision via expensive annotations. In computer vision, Masked Image Modeling (MIM) [12, 79, 64] has emerged as one of the prevalent SSL pre-training paradigms, enabling an efficient pretraining of large models on unlabeled data by reconstructing masked parts of the input images.
footnotetext: Published as a conference paper at ICLR 2025
@@
Figure 2: Left: Downstream evaluations of pre-trained SSL models. MIM-Refiner effectively combines their respective advantages of MIM and ID without suffering from their respective disadvantages. Right: Representation quality of SSL methods evaluated via linear probing. MIM-Refiner advances the state-of-the-art on ImageNet-1K pre-training in low- and high-compute regimes.
In this work, we show that MIM models have different types of blocks: those that mainly improve the pre-training objective and others that are responsible for abstraction within the MIM encoder. The origin of this behavior can be traced back to the fact that MIM architectures usually comprise a large ViT encoder together with a very light-weight decoder. For larger models, the light-weight decoder reaches a point, where it cannot further improve the pre-training objective on its own and passes part of the reconstruction task back to the last encoder blocks. Consequently, the feature quality for downstream tasks of the later blocks degrades, and, somewhat unusual, the representation quality peaks in the middle blocks of the encoder.
@@
Experimentally, we show that within few epochs, MIM-Refiner refines the features of a MIM model to (i) incorporate the beneficial properties of ID objectives (ii) preserves the advantages of the MIM model (iii) exploits the synergies of both methods to improve upon each individual pre-training objective, advancing the state-of-the-art across various benchmarks, see Figure 2. Extensive evaluations show the potential of MIM-Refiner for training large-scale vision foundation models.
@@
    We show via a detailed analysis that MIM models have different types of blocks: those that mainly improve the pre-training objective and others that are responsible for abstraction.
@@
    We experimentally show the effectiveness and generality of MIM-Refiner by refining a multitude of MIM models of various scales, which achieve new state-of-the-art results in a broad range of downstream tasks.
2 Block Types in Masked Image Modeling
@@
Figure 3: Our analysis reveals different block types in MIM models. (a) MIM models are asymmetrically designed where the encoder has most of the parameters, as indicated by the percentages in the bars. (b) The representation quality of MAE encoders (measured by $k$-NN accuracy) peaks in the middle blocks before degrading when the later blocks take over parts of the decoder’s task. Various MIM models and also a semantic segmentation task follow this pattern (see Appendix B.4) (c) The decline in representation quality in later blocks primarily contributes to the degradation in downstream performance. ID methods (represented by iBOT) and our MAE-Refined do not suffer from this issue. (d) Correlation of the relative improvement of reconstruction loss and $k$-NN accuracy per block. The relative improvement is the difference between subsequent blocks divided by the maximum difference over all blocks. Figure (b) and Figure 11 in the appendix show the raw $k$-NN accuracies and reconstruction losses from which the relative improvement is obtained. Middle blocks form abstract representations (large improvements in the $k$-NN accuracy, almost no improvement in reconstruction loss), later blocks take over parts of the reconstruction task (decrease in the $k$-NN accuracy, large improvement in the reconstruction loss). Additional details can be found in Appendix D.2.
We start with an analysis, which – to the best of our knowledge – is the first that clearly reveals the different block types in MIM models. Motivated by these findings, we introduce a targeted refinement process of blocks that harm downstream performance in Section 3.
Different blocks due to asymmetric encoder-decoder design.
MIM models, such as MAE [31] and data2vec 2.0 [6] enable an efficient pre-training of large models. In terms of architecture, the encoder and decoder are intentionally designed asymmetrically. The encoder, on the one hand, is a large ViT that discards 75% of the input patches through masking to drastically reduce training costs. The decoder, on the other hand, operates on the full sequence length – by concatenating mask tokens to the encoded visible patches – and, thus, is typically very lightweight to compensate for the increased number of tokens (Figure 3(a)). As models increase in size, the decoder eventually reaches a point where it cannot further improve the pre-training objective on its own. Consequently, it begins to delegate a portion of the reconstruction task back to the last encoder blocks. This transfer adversely affects the feature quality for downstream tasks associated with those blocks (Figure 3(b)), especially when only few labels are available (Figure 3(c)). We observe this phenomenon by correlating the relative improvement in the reconstruction loss vs the $k$-NN accuracy (Figure 3(d)). Roughly speaking, the blocks of the encoder operate in three different regimes:
@@
What to learn from these findings?
Naïvely using the last encoder block features uses features suited for reconstruction, not for downstream tasks. If lots of labels are available, this can be compensated by fine-tuning the last encoder blocks on the downstream task. However, if not enough labels are available, or the last encoder blocks are not fine-tuned, downstream performance suffers.
One would think that simply using a larger decoder solves these problems. However, there are multiple problems with this solution. (i) The decoder commonly operates on the full sequence length, making it costly to increase its size. (ii) Scaling the decoder can decrease performance as shown, for example, in MAE [31] (Table 1). (iii) Models that can use a deeper decoder (such as CrossMAE [28]) also show degrading representation quality in later blocks (see Appendix Figure 8(d)). Instead of changing established MIM pre-training procedures, we ask the question: can we leverage and further improve the strong intermediate representations of MIM models?
@@ 3 Method
Inspired by NN based contrastive learning [25, 4], we propose Nearest Neighbor Alignment (NNA). NN contrastive objectives introduce an inter-sample correlation by retrieving NNs of samples in a batch and subsequently applying an objective between the samples and their NNs. In practice, the NN retrieval is typically done by tracking samples from previous iterations in a first-in-first-out queue [25, 32]. Therefore, the samples in the queue are from a previous model state which creates a tradeoff between the benefit of the NN-swap and the worse signal from the out-of-sync samples. We argue that the NN-swap does not offer a benefit for negative samples, since they are already different images, and instead, degrades the signal from the contrastive objective. We therefore propose to use the NN only for the alignment of positive samples, as visualized on the right side of Figure 4. Omitting the NN-swap for the negatives stabilizes training for large models and slightly improves performance (see Table 1). NNA is a variant of NNCLR and we visualize their difference in Figure 15.
$\mathcal{Q}$ the NNA objective is:
@@ 4 Experiments
We refine a series of MIM models, namely MAE [31, 72], data2vec 2.0 [6] (abbreviated as D2V2), dBOT [55] and CrossMAE [28]. These models were pre-trained on ImageNet-1K [21], which we also use for the refinement. Models are refined for 20 or 30 epochs using a peak learning rate of 4e-4 with a layer-wise decay of 0.65. ID heads are attached after each block in the last third. Further implementation details are listed in Appendix C. For visual clarity, we compare our best model against previous state-of-the-art models and show the full result tables in the Appendix. We evaluate our models in classification (low- and many-shot), feature evaluation, clustering, transfer learning and semantic segmentation. Various additional evaluations are presented in Appendix B.
@@ 4.1 Ablation Study
Table 1: Ablation study by refining data2vec 2.0 [6] models. Default settings
(a) Adding multiple heads at intermediate features improves $k$-NN accuracy by 0.8%. The best settings include the blocks with the best representation quality (see Figure 3(b)). Adding additional heads before the best blocks does not improve performance while increasing training costs. (b) The benefit doubles to 1.6% for a ViT-H since deeper models degrade more (see Figure 3(b)). Additionally, we investigate where to attach the ID heads in Appendix B.2 where we find that simply attaching ID head to the last 8 (or 12 for ViT-H) blocks to perform best across model scales.
(c) Scheduling the loss weight of each intermediate head is not necessary, i.e.  simply summing all losses is sufficient to achieve good performances. “Uniform Decay” decays the loss weight of all intermediate heads during training. “Staggered Decay” starts by decaying the first head and gradually starts to decay more and more heads as training progresses. “Staggered Step” disables the first head after some time followed by gradually disabling more and more heads. “One Hot” trains the encoder with only one intermediate head at a time where training starts with the earliest intermediate head and gradually iterates through the heads. Details to the loss schedules are in Appendix D.14.
(d) Using the NN swap only for aligning the positive with the anchor gives a small but consistent improvement on smaller models. (e) The improved signal quality is cruicial to avoid training instabilities in larger models. These instabilities manifest in a sudden representation quality drop mid-training leading to a much worse final model.
(f) Relying only on the data-driven augmentation of the NN-swap by omitting color/blur augmentations, is beneficial for certain downstream tasks such as extreme low-shot classification [50]. We show more results without color/blur augmentations in Appendix B.5.
@@ 4.2 Low-shot and Feature Evaluations
We evaluate the ability of our models to perform low-shot classification in Table 2. Additionally, linear probing and $k$-NN accuracy are reported which are computed from the features of the frozen encoder. These metrics are typically correlated to low-shot performance as it indicates that the representation is already linear separable which eases drawing decision boundaries given only few labels. For linear probing, we use the protocol of DINOv2 [62] which includes using features of the last four blocks in its hyperparameter grid. Therefore, linear probing evaluates the features at the end of the encoder, while $k$-NN accuracy evaluates only the features of the last block.
MIM-Refiner drastically improves upon MIM models and other SSL models. In the 1-shot settings D2V2-Refined-H sets a new state-of-the-art of 64.2%, outperforming the 63.6% of MAWS-6.5B [72] which is pre-trained on a private dataset with 2000x the size of ImageNet-1K.
@@ 4.3 Cluster Evaluations
Table 3 shows that MIM-Refiner greatly improves various clustering metrics. The reached ACC of 67.4% outperforms the current state-of-the-art of 61.6% reached by TEMI MSN [1]. Figure 6 illustrate this drastic increase in cluster performance and class separation visually. Additionally, we evaluate combining multiple models using the TURTLE [29] framework in Appendix Table 13.
@@ 4.4 Transfer Learning Evaluations
Table 4: Transferring MIM-Refiner models to other datasets. “VTAB-6” reports the average accuracy over six datasets from the VTAB benchmark [95] and “VTAB-1K” is the average over all 19 datasets of the VTAB-1K benchmark. ADE20K reports the mean intersection over union (mIoU) of a semantic segmentation probe. MIM-Refiner learns general features that can easily be transferred to various datasets and tasks. Table 15 confirms this finding on additional models and Appendix B.11 shows individual VTAB performances for VTAB-6 probing and VTAB-1K fine-tuning.
@@ 4.5 Fine-tuning with Large Amounts of Labels
MIM models typically outperform ID methods when given enough labels to fine-tune the model. As our refinement process employs an ID objective, we investigate whether MIM-Refiner involuntarily degrades performance given an abundance of labels. To this end, we fine-tune MIM models and their refined version on ImageNet-1K [21], iNat18 [38] and ADE20K [97] using a linear classification or UperNet [89] segmentation head. Table 5 shows a small but consistent improvement of MIM-Refiner models.
@@
Model IN-1K Robustness iNat18 ADE20K IN-1K Robustness iNat18
D2V2 86.6 60.2 81.0 54.4 86.6 63.2 79.6
D2V2-Ref. 86.7 60.5 81.6 54.4 86.8 64.1 79.8
dBOT 85.8 55.3 81.9 53.1 87.1 63.7 84.1
dBOT-Ref. 85.9 55.3 82.1 53.3 87.1 64.0 84.5
iBOT 84.8 47.7 76.9 51.1 - - -
Mugs 85.2 46.4 76.9 50.2 - - -
I-JEPA - - - - 84.9 50.9 75.9
Table 5: Full fine-tuning using 100% of the labels. MIM-Refiner consistently improves performance slightly even though ID methods perform worse than MIM models on this benchmark. Appendix Table 18 confirms this pattern on additional MIM models. “Robustness” shows the average performance of the trained IN-1K classifier on robustness datasets (individual results in Table 19). ID and JEPA [3] models (shown in gray) are not competitive with MIM(-Refiner) models.
While improvements in fine-tuning with large amounts of labels might seem marginal when compared to the gains of MIM-Refiner on other benchmarks, it is important to consider that these benchmarks are extremely competitive and advancements thereon are made in small increments. Gains of one percentage point or larger are unrealistic when keeping data and model size constant. Additionally, MIM-Refiner is designed to be compute efficient, requiring only a couple of epochs of training, whereas training from scratch requires multiple hundred epochs. The fact that MIM-Refiner still improves fine-tuning with large amounts of labels slightly shows that our method can efficiently learn strong semantic representations, even improving upon MIM models in their “strong suit” (fine-tuning with large amounts of labels).
4.6 Discussion & Future Work
We experimentally show that MIM-Refiner efficiently combines the advantages of MIM and ID methods without suffering from their respective disadvantages. When comparing to the current state-of-the-art models in ID (iBOT, Mugs), MIM-Refiner consistently outperforms them across a variety of benchmarks, most of the time by large margins. Additionally, the efficiency of MIM-Refiner allows us to scale up model size beyond the largest ID models, without requiring unreasonable amounts of compute. To put it into perspective, the currently largest ID model trained on ImageNet-1K are ViT-L models (300M parameters) whereas MIM-Refiner can effortlessly scale up to 2B parameter models. Even DINOv2 [62], a vision foundation model trained on a private high-quality dataset with 142M images, trains only a 1B parameter model. When comparing against the current state-of-the-art MIM models, MIM-Refiner drastically improves performance in few-shot settings while also showing slight improvements in many-shot settings. Overall, we have not found a single setting where refining MIM models would be undesirable.
MIM-Refiner shows strong scaling behavior, however we are only able to show results on ImageNet-1K due to resource constraints. When scaling up to ImageNet-21K, we would first need to pre-train large-scale MIM models ourselves, including tuning hyperparameters (learning rate, training duration, …). This requires a lot of compute, which is not within our current compute budget. Note that for ImageNet-1K, we simply download the pre-trained checkpoints and only need compute for the refinement process. Additionally, going beyond ImageNet-21K to web-scale datasets is also heavily restricted by the lack of publicly available high-quality datasets for ID training (such as LVD-142M). However, MIM and ID models have been shown to scale well to larger datasets [72, 62], which suggests that also MIM-Refiner would scale well to larger datasets.
5 Related Work
5.1 Pre-training in Computer Vision
5.2 Combining MIM and ID